OpenStudy (anonymous):

help please and show work. z^2-8z=-12

OpenStudy (jdoe0001):

hmm can you factor that?

OpenStudy (jdoe0001):

\(\large z^2-8z=-12\implies \begin{array}{cccllll} z^2&-8z&+12=0\\ &\uparrow &\uparrow \\ &-6-2&-6\cdot -2 \end{array}\)

OpenStudy (jdoe0001):

what do you think?

OpenStudy (anonymous):

thats correct i finally found the answer online. can you help me with more or explain how to do it? i dont understand it.

OpenStudy (anonymous):

@jdoe0001